OpenStudy (anonymous):

f(x)= sqrt x+4 What is f(x+h) + f(x)?

myininaya (myininaya):

f(x+h)=sqrt(x+h)+4 so f(x+h)+f(x)=...........

OpenStudy (yuki):

its sqrt(x+4+h) + sqrt(x) you cannot simplify anymore, but this is usually used to find the derivative of f(x).

myininaya (myininaya):

its sqrt(x+h)+4+sqrt(x)+4=sqrt(x+h)+sqrt(x)+8

myininaya (myininaya):

wait is the sqrt suppose to be over the 4?

myininaya (myininaya):

or just the x?

OpenStudy (anonymous):

Yes. Over it's x+4 all under the squareroot.

myininaya (myininaya):

k so f(x+h)=sqrt(x+h+4) therefore we have f(x+h)+f(x)=sqrt(x+h+4)+sqrt(x+4)

OpenStudy (anonymous):

So it's\[\sqrt{x+h+4}\] + \[\sqrt{x+4}\]

myininaya (myininaya):

yes